OpenStudy (anonymous):

What is the focus of the parabola given by the equation x = 3(y – 1)^2 + 2?

OpenStudy (anonymous):

( y -1)² = (1/3) ( x -2 ) -> 4p = 1/3 => p = 1/12 Thus F ( 25/12, 1)